OpenStudy (anonymous):

Solve the system. y = -1/3 x + 2 and x + 3y = 3 (0, 1) (1, 0) (3, 1/3) (3/2, 0) no solution

OpenStudy (nenadmatematika):

if you plug in y=-1/3x+2 into the other equation you'll get: x+3(-1/3x+2)=3 x-x+6=3 6=3 this is not possible so there's no solution :D